# Decision-aware training of spatiotemporal forecasting models to select a top K subset of sites for intervention

### v2

@@ 4.3 Decision-aware maximum likelihood
[…]
$\textstyle\qquad g_{t}(\phi)$ $=\epsilon+L^{\text{BPR}}(\bm{r}_{t}(\phi),\bm{y}_{t}).$
@@
We call this objective decision-aware ML estimation, or DAML. If the model is well-specified and training data is plentiful, DAML should deliver the same parameters as ML estimation when $\epsilon$ is low enough. However, when the model is misspecified and $\epsilon$ is higher than the BPR delivered by ML-estimated $\phi$, we argue DAML’s constraint will produce better top-K decisions than ML alone, trading lower likelihood for higher BPR. Compared to direct minimization of $\mathcal{J}^{\text{BPR}}$, DAML can deliver similar BPR but more accurate forecasts of $\bm{y}$ for all sites.
To solve in practice, we use the penalty method (Chong & Żak, 2013) to convert to an unconstrained loss:
@@
This DAML formulation makes estimating $\phi$ via gradient descent possible. Here, $\lambda>0$ is a nuisance hyperparameter that must be tuned. When the constraint is not satisfied, larger $\lambda$ values force gradient updates to move parameters further in directions that might satisfy the constraint.
@@ 5 Experiments and Results
Common setup. In each task below, for a fixed task-specific $K$ we first train models for ML and BPR. Using their final BPR values as guidelines, we select a suitable range of $\epsilon$ values for DAML and fit each one, intending to explore intermediate points on the two-objective Pareto frontier (Costa & Lourenço, 2015). When training each objective, we run many random initializations to convergence across a range of learning rates and other hyperparameters. We keep the model $\phi$ from one run that best achieved its objective on a validation set (see App.  C.1 for details), early stopping as needed. A model’s ultimate top-$K$ rankings can have some stochasticity. Thus, later figures show estimated distributions of BPR across 1000 trials of a $M{=}1000$ sample Monte Carlo estimate of $\bm{r}(\phi)$.
@@ 5.1 Synthetic Data
With $L{=}7$ components, the model would be well-specified and could recover the true data-generating process. However, we focus on misspecification, so we fit with $L{=}2$ components. This will hurt likelihood performance, as sites with distinct true means will need to use common means. However, we wish to show that solid top-K decision-making can still happen even with such severe misspecification.
Experiment setup. We use BPR with $K=5$ for this task. For reproducible details, including all hyperparameters, see Supp. D. We only report training set metrics here for simplicity; later tasks assess generalization to test data.
Results and Analysis. From results in Fig. 2, we draw several conclusions. First, training to optimize NLL alone delivers subpar BPR for this task. Second, optimizing for BPR alone yields much better BPR values, suggesting that even this mispecified model can deliver much better top-K decision making than the off-the-shelf ML solution. However, BPR alone yields nonsensical likelihood values, as nothing in the objective forces $\phi$ to be good at modeling the outcomes $y$, only at relative ranking of the $S$ sites. In Fig. 2, we see how our DAML hybrid objective allows a user to traverse the Pareto frontier of likelihood and BPR by enforcing a desired threshold on minimum BPR. As the desired minimum BPR threshold increases, we can sweep the tradeoff between likelihood and BPR. Ultimately, our DAML yields the best high-likelihood, high-BPR solutions in the top-right corner of the Pareto plot.
@@
Figure 3: Pareto frontier of best possible reach (BPR, x-axis) and log likelihood (LL, y-axis) for real-world tasks. Higher is better on both axes. Each panel how the final models estimated by different training methods score on the test set of a forecasting task defined in Sec. 5. To capture the stochasticity of BPR due to our sampling-based ranking estimator, for each model we show an estimated density for BPR over 1000 trials. Across all 3 tasks, our proposed decision-aware ML (DAML) delivers better top-K decisions as measured by BPR than ML estimation. DAML also delivers likelihood comparable to ML methods and much better than directly optimizing BPR.
@@ 5.2 Opioid-related Overdose Forecasting
Motivation. The ongoing opioid overdose epidemic in the United States has incurred roughly 400,000 deaths over the past two decades, with more than 100,000 fatal overdoses after April 2020 (U.S. C.D.C., 2022). Possible evidence-based interventions to mitigate overdose fatalities include the availability of fentanyl testing strips or overdose rescue medications like naloxone. These protective measures work best when directed at small areas that are high-risk, with co-incident education and support for proper follow-through. Public health agencies could use forecasting to help allocate limited resources towards the goal of harm reduction.
@@
Several efforts across the U.S. have developed small-area forecasting models of opioid-related events (Marks et al., 2021b; Neill & Herlands, 2018). A preregistered trial for overdose reduction in Rhode Island (Marshall et al., 2022a) has used BPR-like metrics to evaluate the top-$K$ predictions of models with conventional training. This past work does not rank or train to improve BPR, as we do.
@@
Datasets. We study the capabilities of different methods on two datasets of historical opioid-related overdose mortality. Our IRB provided a Not Human Subjects Research determination for analysis of this decedent data. The first dataset, MA Fatal Overdoses, covers opioid-related overdose deaths in the state of Massachusetts from 2001-2021. This is a private dataset provided by the MA Registry of Vital Records and Statistics. The second dataset, Cook County IL Fatal Overdoses, tracks opioid-involved overdose deaths in the greater Chicago area in 2015-2022. This is an open dataset obtained via the public website of the Cook County Medical Examiner Case Archive (Cook County, IL, 2014-present).
@@
We follow previous evaluations of these datasets in Heuton et al. (2024). Each dataset was processed to a common format providing fatal overdose counts over time and space. For the spatial units, we chose census tracts. Each tract by design contains a mean population of 4000 people, a scale that allows capturing intra-municipality variation in overdoses. For temporal binning, we picked calendar years to reflect the frequency at which health agencies might enact policy changes. Summary facts are in Tab. C.1.
@@
Task. Our forecasting task is to predict the next year’s count of opioid-related fatal overdoses in each census tract. For MA’s $S{=}1620$ tracts, we train on data from 2011-2018, tune hyperparameters on validation data from 2019, and test on 2020 and 2021. For Cook County IL’s $S{=}1328$ tracts, we train on 2015-2019, tune on 2020, and test on 2021 and 2022. These splits follow Heuton et al. (2024).
@@
Given a trained model, we assess heldout likelihood over all sites as well as BPR with $K{=}100$ to measure the model’s where-to-intervene ranking. A $K{=}100$ budget was selected to reflect realistic public health budgets, and is similar to values used in other studies (Marshall et al., 2022a).
@@
Model. A recent benchmark (Heuton et al., 2024) compared many models designed for overdose forecasting, including neural architectures with attention (Ertugrul et al., 2019) and more classical statistical models. A top-performer is negative binomial mixed-effects regression (Marks et al., 2021a). The generative model can be expressed as
@@
Feature vector $\bm{x}_{st}$ includes tract $s$’s overdose gravity, a recent average of opioid-related overdose deaths in tracts spatially near to $s$, as well as measures of sociological vulnerability. See App. C.1.1 for details.
@@
Setup. We followed the common setup described above. For reproducible details specific to overdose tasks, see App. C.1.
Results. Results on test data for both MA and Cook County IL tasks are shown in Pareto frontier plots in Fig. 3. For both datasets, we see that ML training yields suboptimal top-K decisions. Direct optimization of BPR can improve BPR, though gains on test data vary (+0.04 on Cook County; less than +0.01 on MA). Direct BPR solutions do much worse on likelihood, as expected. Our DAML approach provides better top-K decisions than the ML approach, with no visible decay in likelihood.
@@ 5.3 Endangered Bird Forecasting
Motivation. In the 1940s, Whooping Cranes were almost completely extinct in the U.S., with only 20 existing in the wild (Cannon, 1996). Thanks to numerous efforts over the years to preserve their population, roughly 650 cranes live in the wild today. This key species is still listed as endangered due to several threats.
A major flock of cranes, known as the Aransas-Wood population, winters at the Aransas National Wildlife Refuge (ANWR) along the Gulf Coast of Texas (Vartanian, 2023), while spending summers north in Canada. Ecologists wish to actively monitor this population. Regular aerial surveys of the Texas wintering region have been conducted for decades (Taylor et al., 2015). Using binned spatiotemporal data of sighting counts over time from these surveys, we wish to offer data-driven forecasts of where the most cranes will be. This could help stakeholders decide where to send future human monitors or where to place $K$ fixed-location cameras to efficiently track population health.
Data. We use raw data from Taylor et al. (2015), which digitized decades of aerial surveys of ANWR that marked individual crane sightings on paper maps. We processed this ANWR TX Cranes data into a common format of sighting counts over time and space, selecting units to support our goal of using the top-$K$ sites to improve monitoring. For spatial units, we divided the ANWR into 1338 boxes, each 500 meters square. A quality camera or binoculars can reasonably capture a 1.5-meter tall whooping crane in a 250-meter radius. For temporal binning, we use 2 month periods. This choice catches seasonal variation, but avoids how finer scales might burden staff to move cameras too often.
Task. The experiment on this data was trained on years 2002-2006, validated on 2007-2008, and tested on 2009-2010. In this context, a "year" represents one wintering season; winter 2010 means the winter that began in 2010 and ended in early 2011.
We choose $K=50$ for this task. This is an estimate of how many sites scientists might reasonably monitor every two months on a limited budget.
@@
Setup. We followed the common setup described above. See reproducible details in App. C.2.
Results. From results in Fig. 3 (right panel), we find that optimizing NLL or BPR alone will yield poor results in the other metric. In this case, direct optimization of BPR results in the best top-K decisions, but dramatically worse likelihood. Our DAML models improve BPR by up to 0.05 over conventional ML training with modest decay in likelihood.
@@ 6 Discussion
We have addressed two open challenges related to top-K resource allocation problems guided by the best possible reach (BPR) performance metric. We provided a ranking strategy that can outperform simple per-site means. We posed a training objective that strives for high likelihood across all sites while ensuring top-K decisions meet a stakeholder-specified quality level. Our experimental evaluations suggest our approach can better manage tradeoffs in likelihood and BPR than conventional training methods.
There are several limitations to this study. We focused on showing the tradeoffs between likelihood and BPR for a fixed model family in each task without comparing a wide variety of possible models. Our evaluations use fixed $K$ values and do not explore sensitivity to $K$ or other hyperparameters. Practitioners may need multiple metrics to assess overall utility; focusing myopically on BPR may not always be wise. Additionally, our framework assumes any site with large outcome $y_{s}$ is a better candidate for intervention. Future work could explore data-driven site selection that considers how site-specific attributes might make some interventions more or less effective.
@@ Acknowledgments
Authors KH, FSM, and MCH are supported in part by the U.S. National Science Foundation (NSF) via grant IIS # 2338962. Our team also acknowledges support from the American Public Health Association for a data science demonstration project. We are thankful for computing infrastructure support provided by Tufts University, with hardware funded in part by NSF award OAC CC* # 2018149.
@@
Recent work on where-to-intervene decision making has emphasized the importance of incorporating additional constraints into the top-$K$ site selection budget to meet the needs of all constituents. For example, Marshall et al. (2022b) sought a balance of rural and urban sites for intervention in overdose mitigation efforts in the state of Rhode Island. We believe exciting methodological work could extend the ideas in this paper to handle such constraints.
@@
Supplementary Material for
Decision-aware training of spatiotemporal forecasting models
0
Reproducible code used for all experiments is included in the submission’s open-source repository at https://github.com/tufts-ml/decision-aware-topk

### v3

@@ 4.3 Decision-aware maximum likelihood
$\operatorname*{argmin}_{\phi}$- $\sum_{t=1}^{T}\log p_{\phi}(\bm{y}_{t}),\quad\text{s.t.}~g_{t}(\phi)\leq 0~\forall~t,$ (12)
$\text{where}~~\textstyle g_{t}(\phi)=\epsilon+L^{\text{BPR}}(\bm{r}_{t}(\phi),\bm{y}_{t}).$
@@
We call this combined objective decision-aware ML estimation, or DAML. If the model is well-specified and training data are plentiful, DAML should deliver the same parameters as ML estimation when $\epsilon$ is low enough. However, when the model is misspecified and $\epsilon$ is higher than the BPR delivered by ML-estimated $\phi$, we argue DAML’s constraint will produce better top-K decisions than ML alone, trading lower likelihood for higher BPR. Compared to direct minimization of $\mathcal{J}^{\text{BPR}}$, DAML can deliver similar BPR but more accurate forecasts of $\bm{y}$ for all sites. Additionally including the ML objective offers the ability to include an optional prior term $\log{p(\phi)}$ to incorporate any prior knowledge.
To solve in practice, we use the penalty method (chong_introduction_2013) to convert to an unconstrained loss:
@@
This DAML formulation makes estimating $\phi$ via gradient descent possible. Here, $\lambda>0$ is a nuisance hyperparameter that must be tuned. When the constraint is not satisfied, larger $\lambda$ values force gradient updates to move parameters further in directions that might satisfy the constraint. In practice, we set $\lambda$ such that both components of the loss are of similar magnitude during early training.
@@ 5 Experiments and Results
Common setup. In each task below, for a fixed task-specific $K$ we first train models for ML and BPR. Using their final BPR values as guidelines, we select a suitable range of $\epsilon$ values for DAML and fit each one, intending to explore intermediate points on the two-objective Pareto frontier (costa2015ParetoFrontier). When training each objective, we run many random initializations to convergence across a range of learning rates and other hyperparameters (see App.  C.2 for details). We keep the model $\phi$ from one run that best achieved its objective on a validation set, early stopping as needed. A model’s ultimate top-$K$ rankings can have some stochasticity. Thus, later figures show estimated distributions of BPR across 1000 trials of a M=1000M{=}1000 sample Monte Carlo estimate of $\bm{r}(\phi)$ after training is completed.
@@ 5.1 Synthetic Data
With L=7L{=}7 components, the model would be well-specified and could recover the true data-generating process. However, we focus on misspecification, so we fit with L=2L{=}2 components. This will hurt likelihood performance, as sites with distinct true means will need to use common means. However, we wish to show that solid top-K decision-making can still happen even with such severe misspecification.
Experiment setup. We use BPR with $K=5$ for this task. For reproducible details, including all hyperparameters, see App. D. We only report training set metrics here for simplicity; later tasks assess generalization to test data.
Results and analysis. From results in Fig. 2, we draw several conclusions. First, training to optimize NLL alone delivers subpar BPR for this task. Second, optimizing for BPR alone yields much better BPR values, suggesting that even this mispecified model can deliver much better top-K decision making than the off-the-shelf ML solution. However, BPR alone yields nonsensical likelihood values, as nothing in the objective forces $\phi$ to be good at modeling the outcomes $y$, only at relative ranking of the $S$ sites. In Fig. 2, we see how our DAML hybrid objective allows a user to traverse the Pareto frontier of likelihood and BPR by enforcing a desired threshold on minimum BPR. As the desired minimum BPR threshold $\epsilon$ increases, we can sweep the tradeoff between likelihood and BPR. Ultimately, our DAML yields the best high-likelihood, high-BPR solutions in the top-right corner of the Pareto plot.
@@
Figure 3: Pareto frontier of best possible reach (BPR, x-axis) and log likelihood ($L$, y-axis) for real-world tasks. Higher is better on both axes. Each panel how the final models estimated by different training methods score on the test set of a forecasting task defined in Sec. 5. To capture the stochasticity of BPR due to our sampling-based ranking estimator, for each model we show an estimated density for BPR over 1000 trials. Uncertainty in this plot only corresponds to uncertainty in BPR, log likelihood is a point estimate. In all three tasks, our proposed decision-aware ML (DAML) delivers better top-K decisions as measured by BPR than ML estimation. DAML also delivers likelihood comparable to ML methods and much better than directly optimizing BPR. In the Cranes dataset, the DAML objective surprisingly offers better BPR than the BPR-only objective, although the magnitude of this difference is small and perhaps due to the small-scale and sparsity of this dataset.
@@ 5.2 Opioid-related Overdose Forecasting
Motivation. The ongoing opioid overdose epidemic in the United States has incurred over 500,000 deaths in the past decade, with more than 80,000 fatal opioid-related overdoses in 2023 alone (cdc_vrss_2025). Possible evidence-based interventions to mitigate overdose fatalities include overdose education and nalaxone distribution. Scarce resources require local decision makers to allocate these interventions to small areas that are high-risk (allen_provident_2024), with co-incident education and support for proper follow-through. Public health agencies could use forecasting to help allocate limited resources towards the goal of harm reduction.
@@
Several efforts have developed small-area forecasting models of opioid-related events (marksMethodologicalApproachesPrediction2021; neillMachineLearningDrug2018; bauer_small_2023). A preregistered trial for overdose reduction in Rhode Island (marshallPreventingOverdoseUsing2022) used BPR-like metrics to evaluate the top-$K$ predictions of conventionally-trained models. This past work does not rank or train to improve BPR, as we do.
@@
Datasets. We study the capabilities of different methods on two datasets of historical opioid-related overdose mortality. Our IRB provided a Not Human Subjects Research determination for analysis of this decedent data. The first dataset, MA Fatal Overdoses, covers opioid-related overdose deaths in the state of Massachusetts from 2001-2021. This dataset is publicly available upon request from the MA Registry of Vital Records and Statistics. The second dataset, Cook County IL Fatal Overdoses, tracks opioid-involved overdose deaths in the greater Chicago area between 2015 and 2022. This is an open dataset obtained via the public website of the Cook County Medical Examiner Case Archive (cookcountyilMedicalExaminerCase2014).
@@
We follow previous evaluations of these datasets in heutonSpatiotemporalForecastingOpioidrelated2024. Each dataset was processed to a common format of fatal overdose counts per time and spatial unit. For the spatial units, we chose census tracts. Each tract by design contains a mean population of 4000 people, a scale that allows capturing variation in overdoses at the neighborhood level. For temporal binning, we picked calendar years to reflect the frequency at which health agencies might enact policy changes. Summary facts are in Tab. C.1.
@@
Task. Our forecasting task is to predict the next year’s count of opioid-related fatal overdoses in each census tract. For MA’s S=1620S{=}1620 tracts, we train on data from 2011-2018, tune hyperparameters on validation data from 2019, and test on 2020 and 2021. For Cook County IL’s S=1328S{=}1328 tracts, we train on data from 2015-2019, tune on 2020, and test on 2021 and 2022. These splits follow heutonSpatiotemporalForecastingOpioidrelated2024.
@@
Given a trained model, we assess heldout likelihood over all sites as well as BPR with K=100K{=}100 to measure the model’s where-to-intervene ranking. A K=100K{=}100 budget was selected to reflect realistic public health budgets, and is similar to values used in other studies (marshallPreventingOverdoseUsing2022).
@@
Model. A recent benchmark (heutonSpatiotemporalForecastingOpioidrelated2024) compared many models designed for fatal overdose forecasting, including neural architectures with attention (ertugrul_castnet_2019) and more classical statistical models. A top-performing model is negative binomial mixed-effects regression (marks_identifying_2021). The generative model can be expressed as
@@
Feature vector $\bm{x}_{st}$ includes tract $s$’s overdose gravity, a recent average of opioid-related overdose deaths in tracts spatially near to $s$, as well as measures of sociological vulnerability. See App. C.2.1 for details.
@@
Competitor methods. We compare to two other decision-aware methods discussed above: Perturbed Gradient (PG) (huang_decision-focused_2024) and SPO+ (elmachtoubSmartPredictThen2022), as implemented in PyEPO software (tang_pyepo_2023). To be fair, each uses the same model $p_{\phi}$, the ratio estimator to rank sites, and the gradient of this estimator in Eq. (10). We conduct a hyperparameter search over learning rate (and perturbation noise for PG), selecting the model with the best loss value on validation data.
Setup. We followed the common setup described above. For reproducible details specific to overdose tasks, see App. C.2.
Results. Results on test data for both MA and Cook County IL tasks are shown in Pareto frontier plots in Fig. 3. For both datasets, we see that ML training yields suboptimal top-K decisions. Direct optimization of BPR can improve BPR, though gains on test data over ML vary (+0.04 on Cook County; less than +0.01 on MA). Direct BPR and surrogate loss (SPO+, PG) solutions can yield much worse likelihood, as expected. The surrogate loss methods provide lower quality decisions than the DAML and direct BPR approach. Our DAML approach provides better top-K decisions than the ML approach, with no visible decay in likelihood.
@@ 5.3 Endangered Bird Forecasting
Motivation. In the 1940s, whooping cranes were almost completely extinct in the U.S., with only 20 existing in the wild (cannon1996whooping). Thanks to efforts over the years to preserve their population, there are roughly 650 wild cranes today. This key species is still listed as endangered in 2025.
A major flock of cranes, known as the Aransas-Wood population, winters at the Aransas National Wildlife Refuge (ANWR) along the Gulf Coast of Texas (aransas_wood), while spending summers north in Canada. Ecologists wish to actively monitor this population. Regular aerial surveys of the Texas wintering region have been conducted for decades (asurv). Using binned spatiotemporal data of sighting counts over time from these surveys, we wish to offer data-driven forecasts of where cranes may be found. This could help conservationists decide where to send future human monitors or where to place $K$ fixed-location cameras to efficiently track population health.
Data. We use raw data from asurv, which digitized decades of aerial surveys of ANWR that marked individual crane sightings on paper maps. We processed this ANWR TX Cranes data into a common format of sighting counts over time and space, selecting bin sizes to support our goal of using the top-$K$ sites to improve monitoring. Quality cameras or binoculars that could be used to track cranes can reasonably capture a 1.5 meter tall whooping crane in a 250-meter radius. Therefore, for spatial units, we divided the ANWR into 1338 boxes, each 500 meters per side. For temporal binning, we use 2 month periods. This choice catches seasonal variation, but avoids how finer scales might burden staff to move cameras too often.
Task. The experiment on these data was trained on years 2002-2006, validated on 2007-2008, and tested on 2009-2010. In this context, a "year" represents one wintering season; winter 2010 means the winter that began in October 2010 and ended in April 2011. We choose $K=50$ for this task. This is an estimate of how many sites scientists might reasonably monitor every two months on a limited budget.
@@
Setup. We followed the common setup described above. See reproducible details in App. C.3.
Results. From results in Fig. 3 (right panel), we find that optimizing NLL or BPR alone will yield poor results in the other metric. In this case, direct optimization of BPR results in the best top-K decisions, but dramatically worse likelihood. Our DAML models improve BPR by up to 0.05 over conventional ML training with only modest decay in likelihood. Our DAML and direct BPR objectives also deliver better BPR than previous decision-aware methods (PG, SPO+). This is even after providing smarter initializations to these methods, as we found their training had trouble improving on our common random initialization of $\phi$, perhaps due to this task’s much sparser $\bm{y}$ values.
@@ 6 Discussion
We have addressed two open challenges related to top-K resource allocation problems guided by the best possible reach (BPR) performance metric. We provided a ranking strategy that can outperform simple per-site means. We posed a training objective that strives for high likelihood across all sites while ensuring top-K decisions meet a stakeholder-specified quality level. Our experimental evaluations suggest our approach can better manage tradeoffs in likelihood and BPR than conventional training methods or previous decision-aware methods.
There are several limitations to this study. We focused on showing the tradeoffs between likelihood and BPR for a fixed model family in each task without comparing a wide variety of possible models. Only one type of model misspecification is considered each in the synthetic and real-world experiments; perhaps decision-aware objectives are more or less different to maximum likelihood estimates depending on the kind of model misspecification. Our evaluations use fixed $K$ values and do not explore sensitivity to $K$ or other hyperparameters. Practitioners may need multiple metrics to assess overall utility; focusing myopically on BPR may not always be wise. Additionally, our framework assumes any site with large outcome $y_{s}$ is a better candidate for intervention. Future work could explore data-driven site selection that considers how site-specific attributes might make some interventions more or less effective.
@@ Acknowledgments
Authors KH, FSM, and MCH are supported in part by the U.S. National Science Foundation (NSF) via grant IIS # 2338962. Author TJS was supported by the U.S. National Institute on Drug Abuse via grant # R01DA054267. Our team also acknowledges support from the American Public Health Association for a data science demonstration project. We are thankful for computing infrastructure support provided by Tufts University, with hardware funded in part by NSF award OAC CC* # 2018149.
We are grateful to several anonymous reviewers, who pointed us to related work in operations research literature and helped us connect our BPR metric to work on knapsack problems.
@@
Recent work on where-to-intervene decision making has emphasized the importance of incorporating additional constraints into the top-$K$ site selection budget to meet the needs of all constituents. For example, marshallPreventingOverdoseUsing2022 sought a balance of rural and urban sites for intervention in overdose mitigation efforts in the state of Rhode Island. We believe exciting methodological work could extend the ideas in this paper to handle such constraints.
@@
Appendix
This appendix includes additional experimental results and information for understanding and reproducing experiments.
Reproducible code used for all experiments is included in an open-source repository: